OpenStudy (anonymous):

over the set of real numbers, what is the domain of the opposite function g(f(x))?

OpenStudy (anonymous):

\[f(x)=-x-3,g(y)=\sqrt{y} g(f(x))=\sqrt{-3-x}\]

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

would the answer be \[x \le3\]

OpenStudy (anonymous):

you have to solve \[-3-x\geq 0\] which you do in one step by adding \(3\) to both sides

OpenStudy (anonymous):

but you do not get \(x\leq 3\) you get \(x\leq -3\)

OpenStudy (anonymous):

my bad i forgot the sign